### v1

# Learning Dynamics of Meta-Learning in Small Model Pretraining

Abstract
Large language models are powerful but costly. We ask whether meta-learning can make the pretraining of small language models not only better but also more interpretable. We integrate first–order MAML with subset-masked LM pretraining, producing four LLama-style decoder-only models (11M-570M params), and evaluate it on a fundamental NLP task with many settings and real-world applications. Compared with vanilla training, our model (i) reaches the same loss up to 1.6× sooner, (ii) improves F1 on multilingual Universal NER under equal compute, and (iii) makes the training dynamics easy to read: first the network’s representations fan out (“diversify”) and later they collapse into a smaller, shared subspace (“compress”). This two-stage shift shows up as a rise-and-fall in both effective-rank curves and attention-head entropy. The same curves pinpoint which layers specialise earliest and which later reconverge, giving a compact, interpretable signature of meta-adaptation. Code, checkpoints and WandB logs are released.
| pico-maml (Apache 2.0) | |
| pico-maml (Apache 2.0) |
Learning Dynamics of Meta-Learning in Small Model Pretraining
David Demitri Africa††thanks: Corresponding author: dda28@cam.ac.uk Yuval Weiss Paula Buttery Richard Diehl Martinez University of Cambridge
1 Introduction
Small language models (SLMs) are attractive for privacy and energy reasons, but trail large models partly because they converge slowly and plateau early (Godey et al., 2024; Biderman et al., 2023; Diehl Martinez et al., 2024). As opposed to the common method of brute-force scaling, we explore a different axis: learning rules. First-order Model-Agnostic Meta-Learning (MAML) (Finn et al., 2017) promises a learn-to-learn initialization, yet has rarely been applied to decoder models, and its effect on learning dynamics are poorly understood.
We address this by adding meta-learning in model pretraining111Using a lightweight modification of pico-train (Diehl Martinez, 2025), a language model pretraining framework., interleaving ordinary next-token loss (keeps fluency) with 32-way subset-mask (Bansal et al., 2020; Li and Zhang, 2021) episodes (forces rapid binding). Only a tiny MLP head is adapted in the inner loop, so we can track backbone weights without gradient noise. Our contributions are:
1.
Four open SLMs (11M → 570M) trained with this hybrid rule.
2.
A public trainer that logs per-checkpoint singular-value spectra, head entropies and query accuracy to make learning dynamics inspectable.
3.
Evidence that an early "diversify-then-compress" in effective rank predicts final NER F1 improvements.
2 Related Work
Meta-learning for NLP. (MAML; Finn et al., 2017) is an optimisation-based form of meta-learning that learns an initialisation from which a few gradient steps solve new tasks. It has been particularly successful in computer vision classification and reinforcement learning settings (Nichol et al., 2018). Within NLP, MAML has been adapted to a wide spectrum of supervised problems—including text classification, natural language inference, question answering, summarisation and named entity recognition—where a pre-trained encoder such as BERT is further fine-tuned on small datasets (Rajeswaran et al., 2019; Raghu et al., 2021; Hou et al., 2022). These studies therefore operate (i) on encoder-only, masked-language models and (ii) at parameter counts close to the original 110M-parameter BERT. They leave open whether optimisation-based meta-learning helps decoder LMs and whether its benefits persist at larger parameter scales.
Meta-learning for pretraining. Initial NLP attempts applied MAML only at fine-tuning scale (Raghu et al., 2021; Hou et al., 2022). More recent work embeds bilevel objectives directly in pre-training (Miranda et al., 2023; Ke et al., 2021). While promising, these efforts evaluate only a single model size, focus on one downstream task, or release neither code nor weights, limiting reproducibility and obscuring scale trends. We embed meta-learning directly into the pretraining loop, evaluate on various unseen domains in an unseen task, and provide open weights (11M-570M) and layer-wise spectra, filling that gap.
Subset-Mask LMs (SMLMT). SMLMT constructs pseudo-tasks using a subset of vocabulary words (Bansal et al., 2020). Given an unlabeled text corpus, one selects a set of words and builds an -way classification task. For each chosen word, sentences containing it are collected and the word is masked out. The task is then to predict the masked word from the candidates. Li and Zhang (2021) interleaves it with ProtoNet tasks; we interleave with vanilla LM updates and scale to 570M params.
Interpretable training dynamics. Various works discuss the training of language models in phase transitions Olsson et al. (2022); Hoogland et al. (2024), describing broad changes in indicators as the model gains rapidly in capabilities over a short period of time. We study such phase transitions in the context of meta-learning in pretraining.
3 Method
We pretrain four decoder models at 11M, 65M, 181M and 570M parameters with a hybrid objective (Li and Zhang, 2021) that alternates conventional next-token prediction and first-order MAML episodes (Finn et al., 2017). The episodes are generated with Subset-Masked Language Modelling Tasks (SMLMT) (Bansal et al., 2020). This section details the backbone, the meta-learning episode, the optimisation schedule, and the downstream evaluation harness.
3.1 Baselines
The starting point is the open Pico decoder Diehl Martinez (2025), a LLAMA-style Touvron et al. (2023) stack implemented in plain PyTorch. To maintain apples-to-apples comparability with the original models (and as such isolate the effect of introducing MAML to pretraining), we maintain the design choices and hyperparameter choices of the original Pico decoder models. A sequence of decoder blocks receives 2048 input tokens. Each block performs RMSNorm Zhang and Sennrich (2019), grouped-query self-attention Ainslie et al. (2023) with rotary position embeddings Su et al. (2024), and a SwiGLU feed-forward network Shazeer (2020) that expands to before projecting back to the model width . Width is the only scale-dependent hyper-parameter: for the tiny, small, medium and large variants. All models use 12 heads, 4 key–value heads and causal masking.
3.2 Task construction via SMLMT
SMLMT converts unlabelled text into few-shot classification tasks. From the corpus we sample a set of content words, collect sentences that contain each word and replace that word with a single <mask>. The goal is to predict which of the candidates was masked. Each episode supplies support sentences and a disjoint query set. Table 1 shows an episode with city names and supports per class; the query asks the model to complete a new sentence about cherry blossoms. In practice we use and so the task entropy matches the five-bit next-token uncertainty of English text222Shannon’s estimate of printed-English entropy is about 1.3 bits per character (Shannon, 1951); since English BPE tokens span on average about 4 characters (OpenAI, 2025), this implies roughly We therefore use 5 bits per token as a conservative rule of thumb..
| Set | Input (masked) | Label |
| Support ( each) | I visited __ last summer. | Tokyo |
| The sushi festival in __ was unforgettable. | Tokyo | |
| The Big Ben is in __. | London | |
| I caught the tube at __ yesterday. | London | |
| The Seine runs through __. | Paris | |
| She admired the art at the Louvre in __. | Paris | |
| The Forbidden City is in __. | Beijing | |
| I sampled Peking duck in __. | Beijing | |
| Query | I plan to travel to __ to see the cherry blossoms. | Tokyo |
3.3 Optimiser, data, and monitoring
Training runs for 6000 outer updates on four A100 GPUs, with the original Pico-decoder models evaluated at the checkpoint after 6000 steps. Each GPU streams micro batches of 256 sequences from the 30 percent English subset of Dolma Soldaini et al. (2024) that is already tokenised and chunked by Pico Diehl Martinez (2025). The outer optimiser is AdamW with peak learning rate , 2500-step warm-up and cosine decay. Micro batches of 256 sequences are accumulated eight times giving an effective batch of 2048 (1024 for the 11M model). Every 100 steps we evaluate Paloma perplexity (Magnusson et al., 2024) and log the singular values of three attention and three feed-forward matrices to compute effective rank (Diehl Martinez et al., 2024). Query and support accuracies are also tracked.
3.4 Downstream protocol
Named entity recognition (NER), the downstream task for this study, is a fundamental NLP task that identifies and categorizes entities (e.g., persons, organizations, locations) within unstructured text (Chinchor and Robinson, 1997), and is used in healthcare (Kundeti et al., 2016; Polignano et al., 2021; Shafqat et al., 2022), law (Leitner et al., 2019; Au et al., 2022; Naik et al., 2023), business (Putthividhya and Hu, 2011; Alvarado et al., 2015; Zhao et al., 2021), and knowledge graph systems Al-Moslmi et al. (2020). Specifically, we evaluate our models on Universal NER benchmark (Mayhew et al., 2024). UNER v1 comprises three categories of NER evaluation data, each built on top of Universal Dependencies (UD) (Nivre et al., 2016, 2020) tokenization and annotations: publicly available in-language treebanks, parallel UD (PUD) evaluation, and other eval-only sets (Appendix B).
After pretraining we load the checkpoint at step 6000 and attach a fresh linear classifier for UniversalNER. Two fine-tuning settings are used: head-only and full. In the head-only setting the Transformer is frozen so fine-tuning mirrors the inner loop, in the full setting all weights update. Fine-tuning uses AdamW at for at most ten epochs with early stopping on development F1.
4 Model Pretraining
Training-perplexity tradeoff across scales.
The prerequisite for modifying a pretraining method is ensuring the model still learns. All four Pico-MAML variants reach their respective vanilla loss 1.3–1.6× sooner; perplexity improves only in the 11 M model, indicating an optimisation–regularisation trade-off.
| Model | Train Loss @6k | Paloma Perplexity @6k |
| pico-decoder-tiny-1 | 5.31 | 786.85 |
| pico-maml-decoder-tiny-1 | 4.44 | 422.42 |
| pico-decoder-small-1 | 4.14 | 80.25 |
| pico-maml-decoder-small-1 | 3.67 | 113.76 |
| pico-decoder-medium-1 | 3.89 | 77.90 |
| pico-maml-decoder-medium-2 | 3.49 | 78.63 |
| pico-decoder-large-1 | 3.69 | 49.86 |
| pico-maml-decoder-large-1 | 3.49 | 66.62 |
Contrary to expectation, MAML’s inductive bias may favor optimization over regularization. MAML accelerates convergence but degrades out-of-task fluency at medium+ scales.
5 Downstream NER Evaluation
Models are fine-tuned on each dataset in Universal NER (Mayhew et al., 2024; Nivre et al., 2016, 2020) with publicly available train and dev sets 333Namely, ddt, ewt, set, bosque, snk, set, talbanken, gsd, gsdsimp, all. Results (averaged across each finetuning dataset) are shown as micro-F1 scores in Table 3, organized by evaluation group: seen (language with full train/test/dev splits), test-only (using Parallel Universal Dependencies PUD), and test-only low-resource languages (e.g., Cebuano, Tagalog). Heatmaps of individual models are available in Appendix C.3.
| Model | Seen | Test-Only (PUD) | Test-Only (Other) | |||
| Head | Full | Head | Full | Head | Full | |
| tiny (%) | -8.3 | -3.0 | +6.7 | 0.0 | -37.5 | +3.8 |
| small (%) | +2.2 | 0.0 | -17.2 | -0.6 | +46.7 | +7.0 |
| medium (%) | +1.9 | +2.3 | -4.6 | +1.8 | +14.8 | +3.8 |
| large (%) | +6.2 | +4.8 | +7.2 | +3.5 | +2.1 | +8.1 |
The most striking takeaway from this stage is that, when averaged across all evaluation steps in a category, absolute F1 remains low () due to poor zero-shot transfer, especially for logographic scripts. Overall, MAML improves mean F1 by 2-3 points at medium/large scales, confirming a modest “learning-to-learn” effect under full adaptation.444While these results are much worse in comparison to the baseline in the original Universal NER paper (Mayhew et al., 2024), this is likely because is a multilingual model (Conneau et al., 2020) and the pretraining dataset for Pico is entirely in English..
In-language NER gains suggest capacity-dependent meta-learning.
To better understand how meta-initialization influences cross-lingual transfer on seen languages, F1 scores are broken down by dataset within the in-language group. The results are separated by tuning regime to clarify the extent to which meta-learned representations help when only the classifier is updated (head-only) versus when the entire model is fine-tuned.
In the head-only setting (Table 7), absolute F1 scores remain low across most datasets. Tiny models fail to generalize altogether. As seen in MAML shows the strongest and most consistent gains at large scales (Table 4)—most prominently on en_ewt, hr_set, and sv_talbanken-suggesting that episodic pretraining creates more adaptable feature spaces, particularly for common entity types and scripts. On Chinese (zh_gsd, zh_gsdsimp), performance is uniformly poor, confirming the baseline result in Mayhew et al. (2024) that transfer from phonographic to logographic scripts is difficult.
| Model | Danish | English | Croatian | Portuguese | Swedish |
| large (%) | +8.1 | +14.8 | +10.7 | +8.6 | +18.0 |
| Model | Danish | English | Croatian | Portuguese | Swedish |
| tiny (%) | +3.4 | +0.2 | -1.6 | -0.7 | +6.1 |
| small (%) | -3.9 | -4.7 | -1.9 | -2.6 | +4.9 |
| medium (%) | +0.8 | +4.8 | +3.9 | +1.2 | +3.7 |
| large (%) | +3.6 | +4.4 | -0.5 | +4.2 | +2.8 |
In the full setting (Table 5), both vanilla and MAML-pretrained models achieve higher F1 scores across the board. MAML confers consistent +0.01-0.03 gains at medium and large scales, especially for structurally complex languages like Croatian. These relative gains grow as model capacity increases, indicating that larger models benefit more from MAML pretraining. Even in Chinese, where scores are lowest, MAML nudges performance upward. These gains confirm that meta-pretraining does more than support shallow transfer: it reshapes the optimization landscape of the full model in a way that accelerates convergence and improves generalization.
Taken together, these tables validate that MAML pretraining injects a scalable and tunable learning-to-learn signal. However, these average metrics do not tell the full story. Some settings, entity classes, and fine-tuning conditions benefit substantially more than others.
Class-specific prototype bias in entity recognition.
We characterize the specific way MAML pretraining improves performance in NER by breaking down F1 score by entity class in Figure 3.
Meta-pretraining yields a clear capacity threshold in head-only adaptation. Under a frozen backbone, only the large model consistently converts its learned initialization into PER (+0.034) and LOC (+0.023) gains; medium and smaller variants lack the representational bandwidth to rewire person and place distinctions via a shallow classifier. By contrast, even medium and small models see gains in ORG (+0.016 F1) likely because organization names often include distinctive tokens (e.g., “Inc.”, “Corp.”, or “University”) that form rigid, token-level co-occurrence patterns. These simple patterns mirror the pseudo-classification episodes SMLMT generates, so a shallow classifier can latch onto them without requiring deep feature reconfiguration.
Full fine-tuning broadens and amplifies these effects. In the full setting, PER sees the largest MAML-induced lift (up to +0.027 in the large model). LOC improvements (+0.016 at large scale) climb more gradually: place names often span heterogeneous contexts and scripts (e.g. Zagreb vs. Beijing), so meta-pretraining must be supplemented by full gradient flow for location-specific embeddings. ORG continues to enjoy gains (+0.012 at large), reinforcing that organization recognition remains the simplest class to bootstrap from episodic tasks.
Significant zero-shot transfer gains in low-resource languages.
Now, we discuss how inductive biases manifest in zero-shot cross-lingual transfer to low-resource languages—namely, Tagalog (tl) and Cebuano (ceb).
Tagalog and Cebuano are the two most widely spoken native languages in the Philippines, with tens of millions of first-language speakers each. Both are typologically Austronesian and low-resource, but differ significantly. Tagalog is a morphologically rich, predicate-initial language with a complex voice system that encodes syntactic roles (agent, patient, locative, etc.) through verbal affixes and aspect-marking (Kroeger, 1993; Schachter and Otanes, 1983; Ramos, 2021). Word order is flexible and often pragmatically driven, which weakens the utility of positional cues for tasks like named entity recognition. Cebuano is similarly Austronesian but morphologically simpler than Tagalog, with fewer voice alternations and less affixal variation (Tanangkingsing, 2011). It also does not consistently mark syntactic roles with overt case particles; entities must be inferred from context rather than surface markers (Sityar, 2000). Additionally, Cebuano exhibits a distinct orthographic tradition and more conservative vocabulary (e.g., less Spanish borrowing) (Bunye and Yap, 1971), which further distances it from the English-centric token distributions that dominate cross-lingual pretraining datasets. These characteristics make them ideal stress tests for testing the inductive bias of pretraining strategies like MAML.
| Model | Regime | Overall | Cebuano | Tagalog (TRG) | Tagalog (Ugnayan) |
| tiny | head | -100.0% | -100.0% | N/A | N/A |
| small | head | +151.1% | +209.6% | +315.7% | -15.7% |
| medium | head | +24.3% | +16.7% | -20.7% | +534.3% |
| large | head | +9.0% | +0.0% | +57.3% | -37.5% |
| tiny | full | -6.2% | -4.7% | -25.0% | +109.5% |
| small | full | +7.3% | -6.4% | +28.8% | +4.1% |
| medium | full | +0.0% | -1.0% | +1.4% | -2.1% |
| large | full | -8.0% | -14.5% | -1.6% | -0.8% |
In the head-only setting, MAML delivers its greatest impact on small and medium models. For example, the small head jumps from 0.088 to 0.221 overall—an absolute gain of 0.133 F1—and sees particularly large lifts in Cebuano (+0.153) and Tagalog-TRG (+0.262). The medium head also benefits substantially, improving from 0.259 to 0.322. Even the large head picks up a modest +0.030 F1. Only the tiny head collapses, reflecting its inability to form reliable prototypes during meta-training. These patterns suggest that MAML’s episodic learning instills useful, language-agnostic representations in the classifier layers, enabling mid-size heads to generalize token-level cues to new languages without modifying the backbone.
Once we allow full fine-tuning, however, most of MAML’s advantages disappear at higher capacities. The small model retains a small +0.026 F1 edge, but the medium shows no net change and the large actually drops by 0.034. This reversal implies that when every parameter is free to update, the strong gradient signals of full fine-tuning quickly override the meta-learned inductive biases, erasing or even inverting MAML’s earlier head-only gains. The tiny model again underperforms, consistent with its tendency to overfit during meta-training when unconstrained by a fixed backbone.
In the UNER benchmark, Tagalog and Cebuano serve as canonical low-resource, typologically distinct evaluation settings. Overall NER performance remains modest, but, as Table 6 shows, MAML provides meaningful zero-shot boosts in the head-only regime for small and medium models. These gains suggest that even without training exposure to these languages, the inductive biases from English episodic training transfer surprisingly well, at least for token-level prototypes.
6 Learning Dynamics
Despite clear convergence gains, the pretraining metrics alone leave several observations unexplained: the mid-training rebound and double-descent in Paloma perplexity, the abrupt jumps in support versus query accuracy, and the sudden collapse in representation rank. To understand this further, we now turn to a learning-dynamics analysis: tracking episodic support/query performance, classifier head statistics, and proportional effective rank throughout pretraining.
Effective meta-learning has a capacity threshold.
To understand how MAML updates influence learning dynamics during pretraining, we track both support (training set in the inner loop) and query (held out final step in the inner loop) accuracy across training steps (Figure 4).
The small and medium models show clear signs of effective meta-learning. Support accuracy gradually increases and stabilizes around 6–7%, while query accuracy climbs steadily above 40%. This pattern indicates that the models are internalizing a useful task prior, and show smooth convergence with relatively little instability.
The tiny model displays a distinct failure mode. While its support accuracy rises modestly, its query accuracy remains stagnant, hovering just above chance (10%). This suggests the model memorizes support examples but fails to learn task-generalizable features-a canonical symptom of underparameterization in meta-learning (Finn et al., 2017; Rajeswaran et al., 2019). In effect, it lacks the representational bandwidth to encode a shared inductive bias across tasks.
The large model exhibits more complex dynamics. Although it achieves high query accuracy-eventually surpassing 50%-its learning curve is noisier, with sharper fluctuations in both support and query accuracy. These instabilities may arise from the interaction between large-scale parameter updates and stochastic task sampling. While still effective overall, this suggests that large models may require additional stabilization strategies during meta-training (e.g., adaptive inner-loop learning rates or better task normalization).
Interestingly, the large model’s query accuracy (Figure 4) reveals a distinct late-phase jump after 4,500 steps-a grokking-like effect where generalization rapidly improves after a prolonged plateau. This turning point coincides exactly with a stabilization in the head weight variance, suggesting that the model eventually consolidates a useful episodic prior. Such behavior echoes findings in grokking literature (Power et al., 2022; Nanda et al., 2023), where test performance lags training loss for an extended period before suddenly aligning. In the MAML setting, this may correspond to the model first learning how to adapt, before learning to generalize from adaptation.
Taken together, these patterns confirm that meta-learning is most stable within a mid-capacity regime. Models must be large enough to encode reusable structure, but not so large that their learning becomes erratic. These insights help contextualize downstream findings: the best generalization often arises from models that strike a balance between representational power and stable task-level adaptation.
Classifier head weight variance reveals adaptation behavior.
To better understand how episodic adaptation pressures model structure, we track the evolution of classifier head weights across meta-training (Figure 5). Since the head is re-initialized and adapted in every episode, its long-term statistics reflect how the outer loop consolidates across-task regularities.
The top panel shows the standard deviation of head weights. All models exhibit growth in weight variance, indicating increasing expressivity in the task-specific head. The small model diverges most sharply, with its weight variance surpassing all others after 2k steps. This suggests an over-specialization effect: the model learns to adapt aggressively to each task, potentially at the cost of stability.
In the lower panel, the mean of the head weights remains near zero for most models, but the tiny model is an outlier. It accumulates a strong bias in one direction over training, indicating that its head converges toward a fixed mapping that is minimally updated across episodes. This aligns with earlier diagnostics showing that its gradient norms collapse early in training.
These dynamics reinforce the idea that episodic MAML induces a scale-sensitive tradeoff: in higher-capacity models, episodic gradients drive generalizable structure into the shared initialization; in lower-capacity models, this same pressure can cause drift or collapse.
Evidence of representation collapse and reorganization.
To understand how MAML alters internal representations, we track proportional effective rank (PER), a structure-sensitive metric during training applied to both weights and gradients in the attention layers (Figure 33). Following Roy and Vetterli (2007) and Diehl Martinez et al. (2024), effective rank measures the entropy of the singular value spectrum of a matrix, while PER normalizes this by the total dimensionality:
where . PER captures the extent to which the model’s representations or updates span a full-dimensional space; a decline in PER indicates compression or structural specialization.
Across all MAML-pretrained models, a decline in PER over training is observed, indicating that attention representations become progressively lower-rank. But in the large model, this collapse is not smooth. Instead, there is an abrupt, synchronized drop in both PER and Paloma perplexity at roughly step 3000, which follows an earlier rise in perplexity, and yields a “double descent" Belkin et al. (2019) shape (Figure 2). This kind of dynamic-initial generalization, degradation, and then sharp re-convergence-mirrors phenomena seen in grokking and mechanistic phase change literature (Nanda et al., 2023; Power et al., 2022).
We interpret this behavior as a representational phase transition: the model initially fits the objective using diffuse, high-dimensional representations, which are later compressed into task-specialized, low-rank structures. The descent in PER lags behind the initial perplexity gains, and only after this drop does the second descent in Paloma begin. There is no strong evidence of a comparable phase transition in the vanilla models. While the large and medium variants show mild inflection points in loss and perplexity around step 3000, these are gradual and lack the coordinated sharpness seen in the MAML-trained models.
This suggests that MAML’s bilevel updates and episodic task pressure may help reorganize the optimization landscape to favor discrete qualitative shifts in representation. As explored in Olsson et al. (2022); Wang et al. (2024); Hoogland et al. (2024), model training often proceeds in qualitatively distinct stages: from brute-force fitting, to intermediate rule memorization, to compressed algorithmic abstraction. The drop in PER may signal such a transition—from early diffuse representations to compressed heads tuned to solve the repeated structure of SMLMT episodes. This representational transition is also reflected in the model’s adaptation performance. Around the same step where PER and Paloma perplexity undergo a sharp drop (step 3000), both support and query accuracies rise abruptly (see Figure 4). Prior to this point, query accuracy remains relatively flat, indicating that the model struggles to generalize from support to query examples. But after the phase transition, the model rapidly learns to extrapolate, with query accuracy climbing from near random to over 0.5.
This synchrony across metrics—loss, perplexity, support/query accuracy, and PER—provides compelling evidence of a coordinated phase shift in the model’s learning trajectory. When looking into more granular checkpoints (Figure 6), there is clearer evidence that the model transitions from an early stage where it relies on diffuse representations to a later stage where it reorganizes both its representations and update paths into a lower-dimensional, more modular form capable of few-shot generalization. That said, this phase behavior appears scale-sensitive as it is absent in smaller scales. This suggests that the capacity to reorganize may be gated by scale, and that below a certain threshold, the inductive pressure of MAML induces collapse rather than modularization.
7 Conclusion
This study set out to determine whether first–order MAML can make the pretraining of sub-billion–parameter decoder LMs both faster and more intelligible. The evidence gathered across four Pico scales indicates that it can. When the hybrid meta-objective is interleaved with ordinary next-token prediction, every model from 11M to 570M parameters reaches the same cross-entropy loss noticeably sooner than its vanilla counterpart, and the larger two variants carry a two-to-three–point F1 advantage into Universal NER. Further, it shows gains in transfer to low-resource languages, which has the potential to improve the equitability of language technology. Equally important, the spectral logs reveal a striking "diversify-then-compress" shift part-way through training. The moment at which the effective-rank knee appears turns out to be a reliable signal of the final NER score, providing a window into the model’s developing inductive structure that is absent from ordinary loss curves.
Several natural extensions suggest themselves. A first step is to learn whether the same phase transition re-emerges when the corpus is multilingual, which would clarify why cross-script transfer remains the weak point of the present models. Varying which backbone layers adapt, how many steps they receive and how frequently episodes are interleaved may unlock better compute–capability trade-offs. Finally, the clear correlation between the effective-rank collapse and downstream utility hints that spectral diagnostics might serve as a self-supervised early-stopping signal.
Limitations
All training runs stop at exactly six thousand outer steps, a horizon that may be too short for the largest model, so the observed perplexity gap between MAML and vanilla training could shrink or even reverse if optimisation were allowed to continue. Our downstream evaluation focuses on a single task family, sequence labelling, so it remains unclear whether the same advantages would materialise on reasoning or generation-quality benchmarks. Because the corpus is predominantly English, improvements in low-resource or logographic languages remain modest; a more diverse corpus may alter both quantitative and qualitative conclusions. Hyper-parameters such as the hybrid episode probability, the inner-loop learning rate and the 32-way 4-shot episode size were transferred unchanged across scales; dedicated tuning might further modify the trade-off between convergence speed and final perplexity. Models were trained on academic budget, which limited training to 6000 outer steps. Some interesting training dynamics only appear after a very extended period of training, and future work should study this long-term behavior. Finally, each condition was run with a single random seed owing to GPU constraints, so although the phase transition appears robust, the exact magnitude of the gains should be interpreted with caution.
Acknowledgments
This work was supported by a grant from the Accelerate Programme for Scientific Discovery, made possible by a donation from Schmidt Futures. David Demitri Africa is supported by the Cambridge Trust and the Jardine Foundation. Richard Diehl Martinez is supported by the Gates Cambridge Trust (grant OPP1144 from the Bill & Melinda Gates Foundation).
Appendix A Pseudocode
Below is the pseudocode for the MAML and vanilla pretraining setup.
Distributed Subset Masked Language Modeling Tasks (SMLMT) Training
Distributed Autoregressive (AR) Training
A.1 Multi-GPU processing
Pico already uses Lightning-Fabric data parallelism but meta-learning introduces various demands that make multi-GPU processing complicated. A Bernoulli draw is done on one GPU and broadcast so all ranks choose the same objective. Support and query tensors are constructed on rank 0 then scattered, because per-rank random masks would destroy gradient equivalence. Every GPU performs the same ten head updates before any gradient is communicated. A stray early all_reduce would mix gradients from different inner steps, so we place an explicit barrier between inner and outer phases.
Appendix B Universal NER Datasets
To comprehensively evaluate the pretraining method, each permutation of fine-tuning setup ({head-only, full}, fine-tuning dataset ({da_ddt, …, zh_gsdsimp, all}) (where all consists of all available training sets), model size ({tiny, small, medium, large}), and pretraining setup ({vanilla, MAML}) is evaluated, for a total of 160 evaluation runs.
Publicly Available In-language treebanks (9 langs): full train/dev/test splits, identical to the official UD partitions.
da_ddt, en_ewt, hr_set, pt_bosque, sk_snk, sr_set, sv_talbanken, zh_gsd, zh_gsdsimp
Parallel UD (PUD) evaluation (6 langs): single test.txt files, all sentence-aligned across German, English, Portuguese, Russian, Swedish and Chinese.
de_pud, en_pud, pt_pud, ru_pud, sv_pud, zh_pud
Other eval-only sets (3 langs): small test splits for low-resource languages.
ceb_gja (Cebuano), tl_trg (Tagalog TRG), tl_ugnayan (Tagalog Ugnayan)
Appendix C Supplementary Figures
C.1 Supplementary Tables
| Model | da_ddt | en_ewt | hr_set | pt_bosque | sk_snk | sr_set | sv_talbanken | zh_gsd | zh_gsdsimp |
| vanilla_tiny | 0.004 | 0.031 | 0.011 | 0.000 | 0.004 | 0.009 | 0.000 | 0.005 | 0.009 |
| maml_tiny | 0.000 | 0.057 | 0.000 | 0.014 | 0.014 | 0.002 | 0.000 | 0.000 | 0.005 |
| vanilla_small | 0.000 | 0.196 | 0.123 | 0.099 | 0.047 | 0.056 | 0.020 | 0.000 | 0.003 |
| maml_small | 0.004 | 0.156 | 0.162 | 0.104 | 0.063 | 0.044 | 0.000 | 0.003 | 0.005 |
| vanilla_medium | 0.141 | 0.252 | 0.311 | 0.240 | 0.153 | 0.325 | 0.065 | 0.010 | 0.020 |
| maml_medium | 0.087 | 0.288 | 0.329 | 0.243 | 0.136 | 0.362 | 0.108 | 0.005 | 0.010 |
| vanilla_large | 0.247 | 0.366 | 0.401 | 0.337 | 0.178 | 0.422 | 0.261 | 0.034 | 0.039 |
| maml_large | 0.267 | 0.420 | 0.444 | 0.366 | 0.191 | 0.455 | 0.308 | 0.023 | 0.040 |
| Model | da_ddt | en_ewt | hr_set | pt_bosque | sk_snk | sr_set | sv_talbanken | zh_gsd | zh_gsdsimp |
| Large (%) | +8.1 | +14.8 | +10.7 | +8.6 | +7.3 | +7.8 | +18.0 | -32.4 | +2.6 |
| Model | da_ddt | en_ewt | hr_set | pt_bosque | sk_snk | sr_set | sv_talbanken | zh_gsd | zh_gsdsimp |
| tiny (%) | +3.4 | +0.2 | -1.6 | -0.7 | -2.4 | +1.5 | +6.1 | -9.2 | -2.7 |
| small (%) | -3.9 | -4.7 | -1.9 | -2.6 | +3.4 | +0.9 | +4.9 | +1.6 | +4.9 |
| medium (%) | +0.8 | +4.8 | +3.9 | +1.2 | +0.3 | -0.3 | +3.7 | +5.0 | +8.2 |
| large (%) | +3.6 | +4.4 | -0.5 | +4.2 | +5.7 | +1.3 | +2.8 | +3.4 | +5.0 |
C.2 Pretraining Results
We present the unedited pretraining indicators for each pico-maml-decoder model below, as logged on WandB.
C.3 Downstream Evaluation
I present the full downstream evaluation results below, ordered by fine-tuning regime, pretraining setup, and model size.
C.4 Learning Dynamics
We present the learning dynamics indicators for each pico-maml-decoder model below, as logged on WandB.
Appendix D Default pico-maml-train Configurations
| Category | Parameter | Default Value |
| Model | Model Type | pico_decoder |
| Hidden Dimension () | 768 | |
| Number of Layers () | 12 | |
| Vocabulary Size | 50,304 | |
| Sequence Length | 2,048 | |
| Attention Heads | 12 | |
| Key/Value Heads | 4 | |
| Activation Hidden Dim | 3,072 | |
| Normalization Epsilon | ||
| Positional Embedding Theta | 10,000.0 | |
| Training | Optimizer | AdamW |
| Learning Rate | ||
| LR Scheduler | Linear w/ Warmup | |
| Warmup Steps | 2,500 | |
| Gradient Accumulation Steps | 128 | |
| Max Training Steps | 200,000 | |
| Precision | BF16 Mixed | |
| Data | Dataset Name | pico-lm/pretokenized-dolma |
| Batch Size | 1,024 | |
| Tokenizer | allenai/OLMo-7B-0724-hf | |
| Checkpointing | Auto Resume | True |
| Save Every N Steps | 100 | |
| Learning Dynamics Layers |
"attention.v_proj",
"attention.o_proj", "swiglu.w_2" |
| Learning Dynamics Eval Data | pico-lm/pretokenized-paloma-tinsy | |
| Evaluation | Metrics | ["paloma"] |
| Paloma Dataset Name | pico-lm/pretokenized-paloma-tinsy | |
| Eval Batch Size | 16 | |
| Monitoring | Logging Level | INFO |
| Log Every N Steps | 100 | |
| Meta-Learning | Enabled | True |
| Hybrid Ratio | 0.5 | |
| Inner Steps () | 10 | |
| Inner Learning Rate | 0.001 | |
| Support Shots () | 4 | |
| Query Ways () | 32 | |
| Classifier Head Layers | 4 | |
| Classifier Head Hidden Dim | 128 | |
| Classifier Head Dropout | 0.1 | |
| Classifier Head Init Method | xavier | |
| Monitoring | Logging Level | INFO |
| Log Every N Steps | 100 |
Appendix E pico-maml-decoder Models Comparison
| Pico-MAML-Decoder Model Comparison | ||||
| Attribute | tiny | small | medium | large |
| Parameter Count | 11M | 65M | 181M | 570M |
| Hidden Dimension () | 96 | 384 | 768 | 1536 |
| Feed-forward Dim | 384 | 1536 | 3072 | 6144 |
| Training Time (6k steps) | 10h | 15h | 16h | 25h |